OpenStudy (anonymous):

proof sum_{0}^{infty}\frac{ 2^n }{ n! }=e^2

OpenStudy (anonymous):

\[\sum_{0}^{\infty}\frac{ 2^n }{ n! }=e^2\]

OpenStudy (anonymous):

our teacher gave us a hint but i'm totally lost.

OpenStudy (anonymous):

she says use the series for e^x and it should take 2 lines to prove it

OpenStudy (ikram002p):

\[e^x=\sum_{n=0}^{\infty}\frac{ x^n }{ n! }\] \[e^2=\sum_{n=0}^{\infty}\frac{ 2^n }{ n! }\]

OpenStudy (anonymous):

is that it? that is all you have to write to make it a proof?

OpenStudy (ikram002p):

yess since it asked to " use the series for e^x"

OpenStudy (anonymous):

ok thank you very much

OpenStudy (ikram002p):

\(\Huge\color{green}{\text {Np ت}}\)